# Understanding and Evaluating Human Preferences for AI Generated Images with Instruction Tuning

### v1

@@ V-B Performance on AIGC IQA Databases
MINT-IQA (Ours) 66.46
V-C Performance on Traditional IQA Databases
To fruther emphasize the effectiveness and generality of our proposed model, we also test it on three traditional large-scale image quality datasets including two technical quality datasets ( KonIQ-10k [23], SPAQ [24] ,) and one aesthetic quality dataset (AVA [25]). From Tables VI-VIII, we first observe that the proposed model achieves the best performance on all three traditional IQA databases. Our method outperforms the current state-of-the-art methods in terms of both SRCC and PLCC. The results indicate that the proposed model has more powerful representation learning abilities in assessing the quality of traditional natural images compared to other handcrafted-based models and deep learning-based models. Although MINT-IQA is designed for assessing the quality of AIGIs, it can also be applied to evaluate the quality real-captured images. It can be observed that our proposed method significantly outperforms other models on the AVA [25] dataset. Note that the number of images in AVA dataset is larger than the other two datasets, so it may be challenging for other models to learn a broadly representative generic feature in such a large-scale dataset. Fortunately, the proposed model can learn the feature representation from a more diverse range of image content in large-scale datasets, which leads to a significant improvement.
V-D Visual Question Answering Acccuracy
TABLE VI: Results on KonIQ-10k dataset. The best performance results are marked in RED and the second-best performance results are marked in BLUE.
Method SRCC PLCC
BRISQUE [32] 0.665 0.681
ILNIQE [70] 0.507 0.523
HOSA [35] 0.671 0.694
WaDIQaM [40] 0.797 0.805
PQR [71] 0.880 0.884
SFA [72] 0.856 0.872
DBCNN [39] 0.875 0.884
MetaIQA [73] 0.850 0.887
HyperNet [69] (25 crops) 0.906 0.917
MUSIQ [60] 0.916 0.928
TreS [61] 0.907 0.924
LIQE [74] 0.919 0.908
Re-IQA [7] 0.914 0.923
MINT-IQA (Ours) 0.927 0.945
TABLE VII: Results on SPAQ dataset. The best performance results are marked in RED and the second-best performance results are marked in BLUE.
@@ V-D Visual Question Answering Acccuracy
MINT-IQA (Ours) 0.927 0.932
TABLE VIII: Results on AVA dataset. The best performance results are marked in RED and the second-best performance results are marked in BLUE.
@@
MINT-IQA (Ours) 0.776 0.783
@@
TABLE IX: VQA accuracy comparison results. The random accuracy is calculated based on the number of options for the single-choice questions. The best results are highlighted in RED, and the second-best results are highlighted in BLUE
$\text{VQA}_{\text{acc}}$ Quality Authenticity Correspondence
Random 25.0 33.3 33.3
InstructBLIP [55] 8.35 9.08 37.4
Llava [82] 28.3 21.7 45.5
MINT-IQA (Ours) 48.5 57.3 59.4
where $\mathbb{I}(a=b)=1$ if $a$ equals $b$, and $\mathbb{I}(a=b)=0$ if $a$ is different from $b$. In this way, we can measure the performance of the answers generated by language models and evaluate the preference-related understanding capability of these models on AIGIs. As shown in Table IX, compared with the random selection accuracy, InstructBLIP [55], Llava [82], our MINT-IQA model achieves better explanation capability for preference-related questions on AIGIs.
TABLE X: SRCC results of cross-dataset verification from the quality perspective. Bold entries indicate the best performers.
@@ V-E Cross-Dataset Evaluation
We further conduct an experiment to test the generalization ability of MINT-IQA in a more challenging cross-dataset setting. Specifically, we train the models on the training set of one database and directly test the generalization ability on the test set of another database. The cross-dataset evaluation experiment is conducted on both the AIGC IQA databases and traditional IQA databases. As shown in Table X, our proposed model MINT-IQA achieves better results compared to latest models TReS [61] for all tasks.
@@ V-F Ablation Study
To validate the contributions of the different modules in MINT-IQA, we further conduct ablation studies. The results are demonstrated in Tables XI-XV. We report the averaged SRCC, PLCC, KRCC scores from three perspectives on our AIGCIQA2023+ database for each ablation experiment.
@@ V-F1 Comparison between Different Model Backbone
To verify the rationality and effectiveness of the Q-Former in the proposed MINT-IQA model, we change the model backbone to CLIP [20] and BLIP [21] for comparison. We directly add our multi-head quality regression to CLIP [20] and BLIP [21], and train them in a similar way. Table XI manifests that our Q-Former based method outperforms conventional CLIP [20] and BLIP [21] as the model backbone.
@@ V-F2 Comparison between Different Loss Functions
To further validate the effectiveness of the adopted loss function, we compare it with the cross-entropy loss. We can observe from Table XII that the L1 loss is more effective in dealing with our regression-based IQA tasks.
@@ V-F3 Comparison between Different Fix Rates
The training efficiency of a model is crucial in practical applications. Generally, models with fewer active parameters may have better training efficiency but worse performance results. Since the image encoder accounts for the largest proportion of parameters in the model, we conduct ablation experiments to fix some layers of the image encoder. From Table XIII we can observe that freezing the whole model leads to the worst performance, while the fix rate of 0.7 leads to the best performance. It is reasonable since the feature extraction model needs to be trained for this specific task, but due to the dataset size limitation, training the whole model may lead to the overfitting problem. It is also beneficial to consider the balance between the performance and computational efficiency.
@@ V-F4 Contribution of LLM Segmentation and the Feedback Reward Part
To verify the effectiveness of the design of the two improvement parts for the performance, we launch further ablation experiments of adding the LLM segmentation module and feedback reward module, and show the results in Table XIV. Using the prompt after LLM Segmentation as the input leads to better performance than raw prompt. The reward feedback part can also further improve the performance of our model. Moreover, the combination of two further improved modules achieves better results compared to only using one part. This could suggest that more granular input processing and dynamic adjustment of evaluation metrics based on explanatory feedback can lead to a more accurate alignment with human preference judgement.
TABLE XI: Ablation study on model backbone.
@@
TABLE XII: Ablation study on loss function.
@@
TABLE XIII: Ablation study on the fix rate of image encoder.
@@
TABLE XIV: Improvements of LLM segmentation and feedback reward.
@@
V-F5 Contribution of zero convolution
@@
V-F6 Contribution of the vision-language instruction tuning strategy

### v2

@@ V-A Experimental Protocol
[…]
Improvement + 1.3%
Some baseline results are evaluated by ourselves. For traditional handcrafted-based models, they are directly evaluated on the corresponding databases. For vision-language pre-training models, we simply load the pre-trained weights for inference. CLIP score and BLIP score are calculated directly as cosine similarity between text and image embeddings. For deep learning-based models, we use the same training and testing split as the previous literature. For our proposed model, the number of regressors in our model is flexible, corresponding to the varying number of evaluation dimensions in the test databases. Since traditional IQA datasbases have only one dimension for assessment, only one regressor is used for score prediction. The models are implemented with PyTorch and trained on a 40GB NVIDIA RTX A6000 GPU with batch size of 8. The initial learning rate is set to 1e-5, and decreased using the cosine annealing strategy. We employ Adam optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$.
V-B Performance on AIGC IQA Databases
We evaluate the performance of our proposed metric, MINT-IQA, on three recently released AIGC IQA databases including AIGCIQA2023 [25], AGIQA-3K [24], and ImageReward [23]. The performance comparison results are demonstrated in Tables II-IV. Handcrafted-based models show poor performance on AIGC IQA databases, indicating that the handcrafted features mainly targeted for natural images are ineffective for evaluating AIGIs.
TABLE V: Comparison of MINT-IQA v.s. state-of-the-art NR-IQA algorithms on traditional synthetically and authentically distorted datasets. The best performance results are marked in RED and the second-best performance results are marked in blue
[…]
TABLE VI: Results on SPAQ dataset. The best performance results are marked in RED and the second-best performance results are marked in blue.
@@ V-B Performance on AIGC IQA Databases
Q-Bench (InternLM) [48] 0.730 0.750
[…]
Improvement + 0.9% + 0.7%
On the other hand, deep learning-based methods achieve relatively better results. However, they are still far away from satisfactory. MINT-IQA attains state-of-the-art performance across all three AIGC IQA datasetbases, substantially outperforming current IQA models. As shown in Table II, our proposed MINT-IQA model achieves the best performance on our AIGCIQA2023+ from all three perspectives, which demonstrates the effectiveness of MINT-IQA in evaluating human preferences from multiple perspectives. Additionally, as demonstrated in Table III, our model outperforms other methods in measuring the quality perception and text-image alignment on AGIQA-3K [24]. This highlights the versatility of our model in understanding human preference for AIGIs in terms of both quality perception perspective and alignment perspective. Our proposed method also shows superior performance on the ImageReward [23] dataset compared to the ImageReward model, as shown in Table IV, even though we change the training data format from a single image and its MOS to a pair of images and their comparisons. This further manifests the generalization ability of our model.
V-C Performance on Traditional IQA Databases
TABLE VII: Results on AVA dataset. The best performance results are marked in RED and the second-best performance results are marked in blue.
@@ V-C Performance on Traditional IQA Databases
[…]
Improvement + 2.0% + 2.6%
To further emphasize the effectiveness and generality of our proposed model, we also test it on eight traditional large-scale image quality datasets including LIVE [30], CSIQ [31], TID2013 [32], and CLIVE [33], KADID [34], KonIQ-10k [44], SPAQ [36] and AVA [37]. From Tables V-VII, we first observe that the proposed model achieves the best performance on all eight traditional IQA databases. Our method outperforms the current state-of-the-art methods in terms of both SRCC and PLCC. The results indicate that the proposed model has more powerful representation learning abilities in assessing the quality of traditional natural images compared to other handcrafted-based models and deep learning-based models. Although MINT-IQA is designed for assessing the quality of AIGIs, it can also be applied to evaluate the quality of real-captured images. It can be observed that our proposed method significantly outperforms other models on the AVA [37] dataset. Note that the number of images in the AVA dataset is larger than the other seven datasets, so it may be challenging for other models to learn a broadly representative generic feature in such a large-scale dataset. Fortunately, the proposed model can learn the feature representation from a more diverse range of image content in large-scale datasets, which leads to a significant improvement.
V-D Visual Question Answering Acccuracy
TABLE VIII: VQA accuracy comparison results. The random accuracy is calculated based on the number of options for the single-choice questions. The best results are highlighted in RED, and the second-best results are highlighted in blue
$\text{VQA}_{\text{acc}}$ Quality Authenticity Correspondence
Random 25.0 33.3 33.3
InstructBLIP [84] 8.35 9.08 37.4
Llava [117] 28.3 21.7 45.5
[…]
Improvement + 20.2% + 24.0% + 13.9%
@@ V-D Visual Question Answering Acccuracy
where $\mathbb{I}(a=b)=1$ if $a$ equals $b$, and $\mathbb{I}(a=b)=0$ if $a$ is different from $b$. In this way, we can measure the performance of the answers generated by language models and evaluate the preference-related understanding capability of these models on AIGIs. As shown in Table VIII, compared with the random selection accuracy, InstructBLIP [84], Llava [117], our MINT-IQA model achieves better explanation capability for preference-related questions on AIGIs.
TABLE IX: SRCC results of cross-dataset verification from the quality perspective. Bold entries indicate the best performers.
@@ V-E Cross-Dataset Evaluation
We further conduct an experiment to test the generalization ability of MINT-IQA in a more challenging cross-dataset setting. Specifically, we train the models on the training set of one database and directly test the generalization ability on the test set of another database. The cross-dataset evaluation experiment is conducted on both the AIGC IQA databases and traditional IQA databases. As shown in Table IX, our proposed model MINT-IQA achieves better results compared to latest models TReS [89] for all tasks.
@@ V-F Ablation Study
To validate the contributions of the different modules in MINT-IQA, we further conduct ablation studies. The results are demonstrated in Tables X-XIV. We report the averaged SRCC, PLCC, KRCC scores from three perspectives on our AIGCIQA2023+ database for each ablation experiment.
@@ V-F1 Comparison between Different Model Backbone
To verify the rationality and effectiveness of the Q-Former in the proposed MINT-IQA model, we change the model backbone to CLIP [28] and BLIP [29] for comparison. We directly add our multi-head quality regression to CLIP [28] and BLIP [29], and train them in a similar way. Table X manifests that our Q-Former based method outperforms conventional CLIP [28] and BLIP [29] as the model backbone.
@@ V-F2 Comparison between Different Loss Functions
To further validate the effectiveness of the adopted loss function, we compare it with the cross-entropy loss. We can observe from Table XI that the L1 loss is more effective in dealing with our regression-based IQA tasks.
@@ V-F3 Comparison between Different Fix Rates
The training efficiency of a model is crucial in practical applications. Generally, models with fewer active parameters may have better training efficiency but worse performance results. Since the image encoder accounts for the largest proportion of parameters in the model, we conduct ablation experiments to fix some layers of the image encoder. From Table XII we can observe that freezing the whole model leads to the worst performance, while the fix rate of 0.7 leads to the best performance. It is reasonable since the feature extraction model needs to be trained for this specific task, but due to the dataset size limitation, training the whole model may lead to the overfitting problem. It is also beneficial to consider the balance between the performance and computational efficiency.
@@ V-F4 Contribution of LLM Segmentation and the Feedback Reward Part
To verify the effectiveness of the design of the two improvement parts for the performance, we launch further ablation experiments of adding the LLM segmentation module and feedback reward module, and show the results in Table XIII. Using the prompt after LLM Segmentation as the input leads to better performance than raw prompt. The reward feedback part can also further improve the performance of our model. Moreover, the combination of two further improved modules achieves better results compared to only using one part. This could suggest that more granular input processing and dynamic adjustment of evaluation metrics based on explanatory feedback can lead to a more accurate alignment with human preference judgement.
TABLE X: Ablation study on model backbone.
@@
TABLE XI: Ablation study on loss function.
@@
TABLE XII: Ablation study on the fix rate of image encoder.
@@
TABLE XIII: Improvements of LLM segmentation and feedback reward.
@@
TABLE XIV: Ablation study on training strategy.
Dimension Training Strategy SRCC PLCC KRCC
Quality MINT-IQA-joint 0.8801 0.8870 0.6841
MINT-IQA-LoRA 0.8841 0.8866 0.6888
MINT-IQA-single 0.8871 0.8948 0.6946
Authenticity MINT-IQA-joint 0.8229 0.8127 0.6223
MINT-IQA-LoRA 0.8286 0.8137 0.6264
MINT-IQA-single 0.8341 0.8260 0.6373
Correspondence MINT-IQA-joint 0.8226 0.8055 0.6231
MINT-IQA-LoRA 0.8154 0.8067 0.6133
MINT-IQA-single 0.8253 0.8117 0.6268
@@
V-F5 Comparison of Different Training Strategies
We explore the impact of the fine-tuning strategy on the training of the MINT-IQA model, as shown in Table XIV. Specifically, we utilize Low-Rank Adaptation (LoRA) setting the rank parameter to $r=16$. LoRA enables efficient model adaptation by adding a limited number of trainable parameters, which reduces computational costs while preserving performance. Our experimental results demonstrate that incorporating LoRA fine-tuning leads to a modest yet statistically significant improvement in the model’s SRCC. We further explore two distinct training methodologies: MINT-IQA-joint and MINT-IQA-single. The joint training approach simultaneously optimizes three regression losses within a single backward pass, resulting in a unified set of weights that allow for the concurrent inference of three quality scores. This method offers computational efficiency by maintaining only one set of weights. In contrast, the sequential training approach, referred to as MINT-IQA-single, independently optimizes each regression loss through separate backward passes, thereby generating three distinct sets of weights tailored to each specific quality score. Although MINT-IQA-single requires three separate training iterations, it achieves higher SRCC values compared to MINT-IQA-joint, indicating superior alignment with human perceptual judgments. This highlights a trade-off between computational efficiency and assessment accuracy, suggesting that while joint training is more resource-efficient, sequential training provides enhanced performance in quality assessment.
V-F6 Contribution of zero convolution
@@
V-F7 Contribution of the vision-language instruction tuning strategy
@@ VI Conclusion
VII Limitation and Future Works
The current MINT-IQA approach primarily depends the training on subjective human ratings to achieve good consistency with human ratings. While this method provides valuable insights into perceived alignment quality, it is time-consuming and may lack scalability. It is valuable to explore the integration of automated detection mechanisms and high-level semantic analysis to assess image-text alignment in future works. By leveraging advanced techniques such as object detection, scene understanding, and semantic consistency checks, the model can achieve more objective and efficient evaluations, reducing dependence on manual annotations. As multi-modal perception becomes increasingly prevalent, many multi-modal generation works have also been proposed. It is deserved to develop novel fusion strategies that can effectively combine visual, auditory, video, mesh, and potentially other sensory data to perform quality assessment for multi-modal generations, enabling the assessment of diverse media types within a unified framework.
VIII Acknowledgement
This work was supported in part by the National Key R&D Program of China under Grant 2021YFE0206700, in part by the National Natural Science Foundation of China under Grants 62401365, 62271312, 62225112, 62132006, and in part by the Shanghai Pujiang Program under Grant 22PJ1407400.